flying with the wind, a plane flew 1000 miles in 5 hours. Flying against the wind, the plane could fly only 500 miles in the same amount of time. Find the speed of the plane in calm air and the speed of the wind.
Answer by checkley77(12844)

D=RT OR R=D/T
R=1000/5
R=200 MPH. WITH THE WIND.
R=500/5
R=100 MPH. AGAINST THE WIND.
(200-100)/2=100/2=50 MPH IS THE WIND SPEED.
200-50=150 MPH. IS THE SPEED OF THE PLANE IN CALM AIR.
100=50=150 MPH. DITTO.